OpenStudy (anonymous):

what is 10 to the -4 power

OpenStudy (anonymous):

So to do\[10^-4\] As for all negative powers, all you have to do is move the decimal to the left. In this case, four places to the left.

OpenStudy (anonymous):

0.001

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

0.001 is correct.

OpenStudy (anonymous):

how bout to the -1 power

OpenStudy (anonymous):

Then you just move the decimal one place to the left.

OpenStudy (anonymous):

0.01????

OpenStudy (nowhereman):

Maybe you should look at the definition, let x and y be non-negative, real numbers, then \[x^{-y} := \frac{1}{x^y}\]

OpenStudy (anonymous):

10.0, move the decimal one place to the left, 1.00

OpenStudy (anonymous):

And if you wanted to use the definition, which is great, 10^-1 = 1/10^1

OpenStudy (anonymous):

I am sorry, please disregard my 1.00. I made a dumb mistake.

OpenStudy (anonymous):

1/10 = one tenth = .1